Suppose a coin is tossed two times. Let T be the random variable representing the number of tails that occur. Find the values of the random variable T and construct its probability distributions.

1
Expert's answer
2022-03-03T17:19:53-0500

The values of the random variable T: 0, 1, 2.

"P(T=0)=(\frac{1}{2})^2=\frac{1}{4}."

"P(T=1)=(\frac{1}{2})^2*2=\frac{1}{2}."

"P(T=2)=(\frac{1}{2})^2=\frac{1}{4}."